Airline passengers arrive randomly and
independently at the passenger-screening facility at Sea-Tac Airport. The mean arrival rate is 10
passengers per minute. How is this experiment distributed?
a. Poisson
b. Binomial
c. Exponential
d. It does not have a specific form
24. What is the lambda (λ) for this distribution?
a. 5
b. 10
c. 15
d. 20
25. What is the probability of no arrivals in a one-minute period?
a. 0.0000454
b. 0.0006278
c. 0.0072816
d. 0.0123785
26. What is the probability that fewer than 4 passengers arrive in a one-minute period?
a. 0.0889
b. 0.0613
c. 0.0532
d. 0.0103
27. If we look at the number of passengers in a 15 second period, what is the lambda (λ) now?
a. 10
b. 7.5
c. 5
d. 2.5
28. What is the probability that no passengers arrive in a 15 second period?
a. 0.0415
b. 0.0662
c. 0.0821
d. 0.1013
29. What is the probability of at least one passenger arrives in a 15 second period?
a. 0.3562
b. 0.5627
c. 0.7921
d. 0.9179
Answer #1

1)

a. Poisson

24) λ =    10
25) P ( X =    0   ) = e^-10*10^0/0!=   0.0000454

26) P(X<3) = P(X=0) + P(X=1) + P(x=2) + P(X=3) = 0.0103

27) λ = 10/60*15 = 2.5

28) P ( X =    0   ) = e^-2.5*2.5^0/0!=   0.0821

29) P(X≥1) = 1 - P(X=0) = 1-0.0821 = 0.9179